# Borel Homomorphisms from Forests to Kneser Graphs

### v1

@@ 1. Introduction
For every $r\in\mathbb{N}$ there exists $C_{r}\in\mathbb{N}$ such that if a finite $r$-regular graph $H$ has girth at least $r$, then there is no fiid homomorphism from $T_{3}$ to $H$.
@@ 2. Preliminaries
2.2. Kneser graphs and fractional colorings
@@ 2.3. Chromatic numbers of Kneser graphs
Let $n,k\in\mathbb{N}$ with $n\geq 2k$. The Schrijver graph $K^{\prime}(n,k)$ is the induced subgraph of $K(n,k)$ on the set of sets $A\in\binom{n}{k}$ with the property that no two elements of $A$ have a difference of 1 mod $n$.
@@
In general, $|V(K^{\prime}(n,k))|=\frac{n}{k}\binom{n-k-1}{k-1}$. This is well known, but for instance, one can specify a vertex $A$ by choosing a “starting element” $x\in n$, then writing the sequence in $\mathbb{N}^{k-1}$ of lengths of gaps between each pair of consecutive elements of $A$ starting from $x$. Each gap must have positive length and the lengths must add to at most nk1n-k-1, so stars and bars show that there are $\binom{n-k-1}{k-1}$ such sequences. Finally, any element of $A$ can be the starting element, so we have overcounted by a factor of $k$.
@@
Fix $d$ and $g>0$. Let $o\in V(T_{d})$ and $T=T_{d}\upharpoonright B(o,g)$. Let $N=d^{2g}$. $H$ will a graph whose vertex set is the set of injective functions $B(o,g)\to N$ modulo automorphisms of $T$. We think of these as labelings of $T$. Two such labelings $f_{0}$ and $f_{1}$ will be adjacent iff there is a labeing $F:V(T)\to N$ and adjacent vertices $(v_{0},v_{1})\in E(T_{d})$ so that for each $i\in 2$, $(T_{d}\upharpoonright B(v_{i},g),F\upharpoonright B(v_{i},g))\cong(T,f_{i})$. There is an alternate description of the edge set which may be helpful: Identifying $N$ with some set of $N$ points in $2^{\omega}$, call a lift of $f\in H_{i}$ a labeling $x\in F(T_{d})$ which extends $f$. Then $f_{0},f_{1}\in V(H)$ are adjacent iff they have lifts which are adjacent in $S(T_{d})$.
@@
Let $G^{\prime}=G^{\leq 2g}$ denote the Borel graph with $V(G^{\prime})=V(G)$ and where two vertices are adjacent if their distance in $G$ is at most $2g$. Note that $G^{\prime}$ is $d(d-1)^{2g-1}$-regular. Since this is strictly less than $N$, it follows from [13] $G^{\prime}$ admits a Borel proper $N$-coloring, call it $f$.
@@
    Let $A\subseteq V(\mathcal{H})$ be Baire measurable and nonmeager. For each $i\in d$ there exist $x,x^{\prime}\in A$ such that $(x,x^{\prime})$ is an edge in $\mathcal{H}$ with label $i$.
@@
We now describe a method of constructing elements of $G$ that yields a natural bijection between $V(G)$ and $V(\mathcal{H})\times\omega^{\mathbb{Z}_{2}^{*d}\setminus\{1\}}$. First, let $x\in V(\mathcal{H})$ and $i\in d$. Let $n_{0}^{i}<n_{1}^{i}<\cdots$ enumerate the elements $n$ of $e^{-1}(i)$ for which $s_{n}\subseteq x$. Then for each $m$, $x$ is adjacent to the real obtained by flipping the $n_{m}^{i}$-th bit of $x$ and the edge to this vertex has color $i$. Furthermore this is enumerates all such edges. Let $f_{i,m}(x)$ denote the neighbor corresponding to $m\in\omega$. We will need the following later.

### v2

@@ 1. Introduction
For every $r\in\mathbb{N}$ there exists $C_{r}\in\mathbb{N}$ such that if a finite $r$-regular graph $H$ has girth at least $C_{r}$, then there is no fiid homomorphism from $T_{3}$ to $H$.
@@ 2. Preliminaries
$2.$ Kneser graphs and fractional colorings
@@ 2.3. Chromatic numbers of Kneser graphs
Let $n,k\in\mathbb{N}$ with $n\geq 2k$. The Schrijver graph $K^{\prime}(n,k)$ is the induced subgraph of $K(n,k)$ on the set of sets $A\in\binom{n}{k}$ with the property that no two elements of AA have a difference of 1 mod $n$.
@@
In general, $|V(K^{\prime}(n,k))|=\frac{n}{k}\binom{n-k-1}{k-1}$. This is well known, but for instance, one can specify a vertex AA by choosing a “starting element” $x\in n$, then writing the sequence in $\mathbb{N}^{k-1}$ of lengths of gaps between each pair of consecutive elements of AA starting from $x$. Each gap must have positive length and the lengths must add to at most nk1n-k-1, so stars and bars show that there are $\binom{n-k-1}{k-1}$ such sequences. Finally, any element of AA can be the starting element, so we have overcounted by a factor of $k$.
@@
Fix $d$ and $g>0$. Let $o\in V(T_{d})$ and $T=T_{d}\upharpoonright B(o,g)$. Let $N=d^{2g}$. $H$ will a graph whose vertex set is the set of injective functions $B(o,g)\to N$ modulo automorphisms of $T$. We think of these as labelings of $T$. Two such labelings $f_{0}$ and $f_{1}$ will be adjacent iff there is a labeing $F:V(T)\to N$ and adjacent vertices $(v_{0},v_{1})\in E(T_{d})$ so that for each $i\in 2$, $(T_{d}\upharpoonright B(v_{i},g),F\upharpoonright B(v_{i},g))\cong(T,f_{i})$. There is an alternate description of the edge set which may be helpful: Identifying NN with some set of NN points in $2^{\omega}$, call a lift of $f\in H_{i}$ a labeling $x\in F(T_{d})$ which extends $f$. Then $f_{0},f_{1}\in V(H)$ are adjacent iff they have lifts which are adjacent in $S(T_{d})$.
@@
Let $G^{\prime}=G^{\leq 2g}$ denote the Borel graph with $V(G^{\prime})=V(G)$ and where two vertices are adjacent if their distance in $G$ is at most $2g$. Note that $G^{\prime}$ is $d(d-1)^{2g-1}$-regular. Since this is strictly less than NN, it follows from [13] $G^{\prime}$ admits a Borel proper NN-coloring, call it $f$.
@@
    Let $A\subseteq V(\mathcal{H})$ be Baire measurable and nonmeager. For each $i\in d$ there exist $x,x^{\prime}\in A$ such that $(x,x^{\prime})$ is an edge in $\mathcal{H}$ with label ii.
@@
We now describe a method of constructing elements of $G$ that yields a natural bijection between $V(G)$ and $V(\mathcal{H})\times\omega^{\mathbb{Z}_{2}^{*d}\setminus\{1\}}$. First, let $x\in V(\mathcal{H})$ and $i\in d$. Let $n_{0}^{i}<n_{1}^{i}<\cdots$ enumerate the elements $n$ of $e^{-1}(i)$ for which $s_{n}\subseteq x$. Then for each $m$, $x$ is adjacent to the real obtained by flipping the $n_{m}^{i}$-th bit of $x$ and the edge to this vertex has color ii. Furthermore this is enumerates all such edges. Let $f_{i,m}(x)$ denote the neighbor corresponding to $m\in\omega$. We will need the following later.